OpenStudy (anonymous):

how do we integrate z/(x^2+z^2) with respect to x the answer is arctan(x/z)

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ z }{ x ^{2}+z ^{2} }dx\]

OpenStudy (anonymous):

this is a formula you use to compute the integral of a function with this structure.

OpenStudy (anonymous):

You will be able to relate your question simply to the standard integral \[\large \frac{1 }{ a^2+x^2 } =\frac{1 }{a } \times \arctan (\frac{x}{ a })\]

OpenStudy (anonymous):

if Left hand side is the function that has to integrate with respect to x

OpenStudy (anonymous):

is 1/a the derivative of x/a

OpenStudy (anonymous):

is tht why its there?

OpenStudy (anonymous):

1/a is part of the standard formula. I'll work it out for you.

OpenStudy (anonymous):

\[z \int\limits_{}^{}\frac{ 1 }{ x ^{2}+z ^{2} }dx \]

OpenStudy (anonymous):

thnk you id like to see why

OpenStudy (anonymous):

\[z*\frac{ 1 }{ z }\arctan(\frac{ x }{ z })\rightarrow \arctan(\frac{ x }{ z })\]

OpenStudy (anonymous):

in this case z is a scalar.

OpenStudy (anonymous):

so typically z would be represented as a number. I think they are using zahlen which denotes integer values.

OpenStudy (anonymous):

why is arc tan (x/z) and where does the 1/z come from??

OpenStudy (anonymous):

1/z is always part of the equation. if you want me to do a proof, I'm afraid I don't have enough time for that but maybe someone else would.

OpenStudy (anonymous):

the formula for most of the integral is coming fromits derivative remember derivation is the opposite of integration and wise versa. can you say what is th derivative of f(x) = arctan x ??

OpenStudy (anonymous):

it is 1/1+x^2

OpenStudy (anonymous):

well,good.. then what will be derivative of arctan(x/a)

OpenStudy (anonymous):

i dunno

OpenStudy (anonymous):

ok.. instead of x it is x/a.. substitute that value in your answer above

OpenStudy (anonymous):

and rember x/a is another function so you need to apply the chain rule of differentiation

OpenStudy (anonymous):

1/(1+(x/a)^2)*(1/a)

OpenStudy (anonymous):

Yup........... You got it.. simplify it

OpenStudy (anonymous):

I mean, I can do the integral if you want it in all the gruesome detail...

OpenStudy (anonymous):

a/(x2+a^2)

OpenStudy (anonymous):

yea shure if its not too complicated

OpenStudy (anonymous):

well you are right ... now compare with your question......you got your answer

OpenStudy (anonymous):

that is you got \[\frac{ d }{ dx }(\arctan(x/a) = \frac{ a }{ x^2+a^2 }\]

OpenStudy (anonymous):

I mean, there isn't much too it though. The formula is basically where it comes from... Now that I think about it, I don't know how much you can actually write. I was running it by wolfram and all they do is a variable substitution and then use that int (1/u^2+1)=arctan(u) :/

OpenStudy (anonymous):

now itegrat on both sides.. you got answer

OpenStudy (anonymous):

Remember Integration is the opposite of differentiation @SABREEN .. now what will be tha answer?

OpenStudy (anonymous):

so than the integral of a/(x^2+a^2) is arctan (x/a)

OpenStudy (anonymous):

thanks=)

OpenStudy (anonymous):

yes.. Welcome.. Now should I explain how 1/a comes ???? ;)

OpenStudy (anonymous):

Next one in line: \[\int\limits \tanh^{-1}(ax)dx\]

OpenStudy (anonymous):

loll nahh i got this bro=)

OpenStudy (anonymous):

@SABREEN ....Good... Welcome and enjoy.... @malevolence19 .......you could solve by following the method i described here...only difference is x = ax and remember ax is a fuction.. :)

OpenStudy (anonymous):

I mean, I know how to to do it lol. It's a integration by parts problem once you look up the derivative to tanh^(-1)(x)